# Integrating Low-Altitude SAR Imaging into UAV Data Backhaul

### v1

@@ II-C SAR Echo Model
We assume that the received OFDM echo are reflected by $Q$ resolvable scatterers/pixels, where the $q$th one is positioned on the grid of $(x_{q},y_{q})$. Then the echo in the TF domain can be formulated as
@@
where $c$, $\bar{\alpha}_{q}$, $R_{q,m}$ and $z_{n,m}$ represent the speed of light, the radar cross section (RCS) coefficient caused by the $q$th scatterer within the radar beam footprint, the instantaneous slant range between the UAV and the $q$th scatterer, and the additive white Gaussian noise, respectively. Note that $\bar{\alpha}_{q}\sim\mathcal{CN}(0,\sigma^{2}_{\alpha_{q}})$ and $z_{n,m}\sim\mathcal{CN}(0,\sigma^{2})$.
@@
where $\bar{R}_{q}=\sqrt{x^{2}_{q}+H^{2}_{p}}$. Note that the approximation holds when the azimuth synthetic aperture length $vT_{a}$ is much smaller than the slant range $\bar{R}_{q}$ and the radar operates under a small squint angle [cumming2005digital]. In LAWNs, the term $vT_{a}$ for the UAV platform can reach on the order of hundreds of meters, while its altitude can be on the order of several thousands of meters. In addition, the small-squint-angle condition is naturally satisfied in broadside mode. Therefore, the approximation in (8) holds.
Then substituting (8) into (7) and rearranging terms, we obtain
@@
where $\alpha_{q}=d_{q}e^{-j\frac{4\pi}{c}f_{c}\bar{R}_{q}}$. In addition to separated range and azimuth terms, the range-cell migration (RCM) term represents the coupling between range and azimuth phases in the TF domain. Therefore, conventional 2D-FFT approach [du2025prob, keskin2024fundamental] cannot be directly utilized, making it difficult to achieve their independent focusing.
@@
$\bm{Y}=\bm{H}\odot\bm{S}+\bm{Z},$ (10)
where $\bm{S}\in\mathbb{C}^{N\times M}$ and $\bm{Z}\in\mathbb{C}^{N\times M}$ represent the random symbol and noise matrices, respectively, and $\bm{H}\in\mathbb{C}^{N\times M}$ denotes the imaging channel matrix, expressed as
$\bm{H}=\sum\nolimits^{Q}_{q=1}\alpha_{q}\underbrace{\bm{b}\bm{c}^{H}\odot\bm{E}}_{\bm{H}_{q}},$ (11)
where $\bm{b}\in\mathbb{C}^{N\times 1}$ with the $n$th element as $\bm{b}]_{n}=e^{-j\frac{4\pi}{c}n\Delta f\bar{R}_{q}}$, $\bm{c}\in\mathbb{C}^{M\times 1}$ with the $m$th element as $\bm{c}]_{m}=e^{-j\frac{4\pi}{c}f_{c}\Delta R_{q,m}}$, and $\bm{E}\in\mathbb{C}^{N\times M}$ with the $(n,m)$th element as $\bm{E}]_{n,m}=e^{-j\frac{4\pi}{c}n\Delta f\Delta R_{q,m}}$.
@@
In OFDM-SAR imaging, the reconstruction problem can be expressed as a linear model by reformulating (10) as
${\bm{y}}=\tilde{\bm{H}}\bm{\alpha}+{\bm{z}},$ (12)
where ${\bm{y}}=\mathrm{vec}(\bm{Y})$, $\tilde{\bm{H}}=\left[\mathrm{vec}(\bm{H}_{1}\odot\bm{S}),\cdots,\mathrm{vec}(\bm{H}_{Q}\odot\bm{S})\right]$, $\bm{\alpha}=[\alpha_{1},\cdots,\alpha_{Q}]^{T}$, and $\bm{z}=\mathrm{vec}(\bm{Z})$. From the perspective of parameter estimation, SAR imaging is equivalent to the estimation of $\bm{\alpha}$ [zhang2014ofdm, zhang2014irci]. Since $\bm{z}$ is Gaussian and satisfies $\mathbb{E}\{\bm{z}\}=\bm{0}$ and $\bm{R}_{z}=\mathbb{E}\{\bm{z}\bm{z}^{H}\}=\sigma^{2}\bm{I}$, the maximum likelihood estimation of $\bm{\alpha}$ is equivalent to its least-squares counterpart as
$\hat{\bm{\alpha}}=\left(\tilde{\bm{H}}^{H}\tilde{\bm{H}}\right)^{-1}\tilde{\bm{H}}^{H}\bm{y}.$ (13)
However, the number of pixels $Q$ is typically enormous. Consequently, $\tilde{\bm{H}}\in\mathbb{C}^{NM\times Q}$ becomes too large to store, also making direct least-squared inversion of a $Q$-dimensional matrix computationally intractable. To address this challenge, SAR commonly exploits the inherent Fourier-like structure of $\tilde{\bm{H}}$, enabling efficient image reconstruction through fast 2D Fourier-based approaches such as RD and Omega-K algorithms [cumming2005digital]. These approaches approximate the least-squared solution with much lower complexity while maintaining near-optimal focusing performance, which is clarified in Sec. III.
@@ III SAR Imaging Approach
In this section, we elaborate on the procedure of integrating SAR imaging into UAV data backhaul. There are two key differences compared to the conventional RD algorithm. First, we propose to exploit several TF-domain filtering schemes to allieviate the impact of data randomness. Second, the azimuth signal is not a strict quadratic-phase signal, since the random data imposes amplitude modulation on it, and the approximation of its Doppler spectrum requires further discussion.
@@ III-A TF Filtering
Different from conventional SAR imaging based on chirp waveforms, SAR using OFDM communication signals may suffer from severe imaging degradation due to the signaling randomness and the employed TF-domain filtering schemes. To mitigate this effect, the first step is to compensate for the discrete symbols drawn from a given QAM constellation.
@@
$\bm{y}=\bm{A}\bm{h}+\bm{z},$ (14)
where $\bm{A}=\mathrm{diag}(\mathrm{vec}(\bm{S}))$, and $\bm{h}=\mathrm{vec}(\bm{H})$ which involves $\bm{\alpha}$. Then we arrive at $\mathbb{E}\{\bm{h}\}=\bm{0}$ and $\bm{R}_{h}=\mathbb{E}\{\bm{h}\bm{h}^{H}\}=\sum^{Q}_{q=1}\sigma^{2}_{\alpha_{q}}\bm{I}$, which can be similarly proved by referring to [keskin2024fundamental, Lemma 1].
@@
By rewriting the model as (14), we shift the focus to the channel (observation) space, where $\bm{h}$ represents the equivalent imaging channel. The diagonal structure of $\bm{A}$ decouples the observations, enabling efficient $NM$ parallel estimation. Notably, from a parameter-space perspective like that in (12), estimating $\bm{h}$ implicitly contains information of $\bm{\alpha}$. Consequently, this formulation provides a natural two-stage estimation strategy: 1) channel-space estimation: estimate $\bm{h}$ independently for each subchannel; 2) parameter-space inversion: recover the physical scatterer parameters $\bm{\alpha}$ from $\bm{h}$ in accordance with their linear relationship as $\bm{h}=\mathrm{vec}(\bm{H})=\mathrm{vec}\left(\sum_{q}\alpha_{q}\bm{H}_{q}\right)=\sum_{q}\alpha_{q}\mathrm{vec}\left(\bm{H}_{q}\right)$, through 2D 2D-DFT-based algorithms.
Our current aim is to estimate $\bm{h}$, formalized as $\hat{\bm{h}}=\bm{G}\bm{y}$, where $\bm{G}$ represents the TF-domain filtering matrix. Inspired by [du2025prob], we now consider three filtering strategies.
@@ III-A1 Reciprocal Filtering (RF)
$\hat{\bm{h}}_{\mathrm{RF}}=\arg\min_{\bm{h}}\ \|\bm{y}-\bm{A}\bm{h}\|^{2}_{F},$ (15)
@@
$\hat{\bm{h}}_{\mathrm{RF}}=\underbrace{\left(\bm{A}^{H}\bm{A}\right)^{-1}\bm{A}^{H}}_{\bm{G}_{\mathrm{RF}}}\bm{y}.$ (16)
Thanks to the diagonal structure of $\bm{A}$, (16) is equivalent to
$\hat{\bm{H}}_{\mathrm{RF}}=\frac{\bm{Y}\odot\bm{S}^{*}}{|\bm{S}|^{2}}=\frac{\bm{Y}}{\bm{S}},$ (17)
where $\hat{\bm{H}}_{\mathrm{RF}}$ is the RF estimation of $\bm{H}$.
@@ III-A2 Matched Filtering (MF)
MF is designed based on the rule of maximum signal-to-noise ratio (SNR) output, which may be quantified as
$\mathrm{SNR}_{\mathrm{out}}=$ $\frac{\mathbb{E}\{\|\bm{G}_{\mathrm{MF}}\bm{A}\bm{h}\|^{2}_{F}\}}{\mathbb{E}\{\|\bm{G}_{\mathrm{MF}}\bm{z}\|^{2}_{F}\}}=\mathrm{SNR}_{\mathrm{in}}\cdot\frac{tr\left(\bm{G}_{\mathrm{MF}}\bm{A}\bm{A}^{H}\bm{G}^{H}_{\mathrm{MF}}\right)}{tr\left(\bm{G}_{\mathrm{MF}}\bm{G}^{H}_{\mathrm{MF}}\right)}$ (18)
$\leq$ $\mathrm{SNR}_{\mathrm{in}}\cdot\frac{tr\left(\bm{A}\bm{A}^{H}\right)tr\left(\bm{G}^{H}_{\mathrm{MF}}\bm{G}_{\mathrm{MF}}\right)}{tr\left(\bm{G}_{\mathrm{MF}}\bm{G}^{H}_{\mathrm{MF}}\right)}$
$=$ $\mathrm{SNR}_{\mathrm{in}}\cdot tr\left(\bm{A}\bm{A}^{H}\right)=NM\cdot\mathrm{SNR}_{\mathrm{in}},$
where $\mathrm{SNR}_{\mathrm{in}}=\sum^{Q}_{q=1}\sigma^{2}_{\alpha_{q}}/\sigma^{2}$, and the maximum $\mathrm{SNR}_{\mathrm{out}}$ is attained when $\bm{G}_{\mathrm{MF}}=\beta\bm{A}^{H}$. Without loss of generality, we use $\beta=1$, leading to
$\hat{\bm{h}}_{\mathrm{MF}}=\bm{A}^{H}\bm{y}.$ (19)
Consequently, the imaging channel matrix is estimated through
$\hat{\bm{H}}_{\mathrm{MF}}=\bm{Y}\odot\bm{S}^{*}.$ (20)
@@ III-A3 Wiener Filtering (WF)
${\bm{G}}_{\mathrm{WF}}=\arg\min_{\bm{G}}\ \mathbb{E}\{\|\bm{h}-\bm{G}\bm{y}\|^{2}_{F}\},$ (21)
@@
$\hat{\bm{h}}_{\mathrm{WF}}=\underbrace{\bm{R}_{h}\bm{A}^{H}\left(\bm{A}\bm{R}_{h}\bm{A}^{H}+\bm{R}_{z}\right)^{-1}}_{\bm{G}_{\mathrm{WF}}}\bm{y}.$ (22)
@@
$\hat{\bm{H}}_{\mathrm{WF}}=\frac{\bm{Y}\odot\bm{S}^{*}}{|\bm{S}|^{2}+\mathrm{SNR}^{-1}_{\mathrm{in}}}.$ (23)
Combining (17), (20) and (23), we summarize the formulation of TF-domain filtering matrix $\bm{G}$ under different filtering schemes as
$\bm{G}=\left\{\begin{array}[]{ll}\frac{1}{\bm{S}},&\mathrm{RF},\\ \bm{S}^{*},&\mathrm{MF},\\ \frac{\bm{S}^{*}}{\left|\bm{S}\right|^{2}+\mathrm{SNR}_{\mathrm{in}}^{-1}},&\mathrm{WF},\end{array}\right$. (24)
@@
$\bm{Y}^{\mathrm{tf}}\overset{\triangle}{=}\hat{\bm{H}}=\bm{Y}\odot\bm{G}=\bm{H}\odot\bm{S}\odot\bm{G}+\bm{Z}\odot\bm{G},$ (26)
@@
$\mathrm{MSE}=$ $\mathbb{E}\left\{\left\|\hat{\bm{H}}-\bm{H}\right\|^{2}_{F}\right\}$ (28)
$=$ $\mathbb{E}\left\{\left\|\bm{H}\odot\bm{S}\odot\bm{G}-\bm{H}+\bm{Z}\odot\bm{G}\right\|^{2}_{F}\right\}$
@@
where the third equality holds due to the facts that: 1) the statistical independence between $\bm{H}$ and $\bm{Z}$, and 2) the statistical characteristics of random variables in (28) are invariant across subcarriers and symbols. Readers are referred to [du2025prob] for a detailed derivation.

### v2

@@ II-C SAR Echo Model
We assume that the received OFDM echo are reflected by $Q$ resolvable scatterers/pixels, where the $q$th one is positioned on the grid of $(x_{q},y_{q})$. In practical low-altitude UAV SAR scenarios, the wireless channel is inherently time-varying due to platform motion, introducing Doppler-induced phase variations across the coherent processing interval. Then the echo in the TF domain can be formulated as [^3] Under typical 5G-NR configurations, the channel dynamics-induced ICI effect in a moving-platform SAR system remains limited. Specifically, the maximum normalized Doppler shift $\eta=\frac{2vf_{c}}{c\Delta f}\approx 0.04$ for $f_{c}=3.5$ GHz, $\Delta f=30$ kHz, and $v=50$ m/s. Notably, this corresponds to the worst case, while in practical side-looking SAR geometries the effective Doppler is much smaller. Therefore, the induced ICI is generally weak as $\eta<0.1$ [34] and can be reasonably neglected in subsequent modeling.
@@
where $c$, $\bar{\alpha}_{q}$, $R_{q,m}$ and $z_{n,m}$ represent the speed of light, the radar cross section (RCS) coefficient caused by the $q$th scatterer within the radar beam footprint, the instantaneous slant range between the UAV and the $q$th scatterer, and the additive white Gaussian noise, respectively. For analytical convenience, it is assumed that $\bar{\alpha}_{q}\sim\mathcal{CN}(0,\sigma^{2}_{\alpha_{q}})$ and $z_{n,m}\sim\mathcal{CN}(0,\sigma^{2})$.
@@
where $\bar{R}_{q}=\sqrt{x^{2}_{q}+H^{2}_{p}}$. Note that the approximation holds when the azimuth synthetic aperture length $vT_{a}$ is much smaller than the slant range $\bar{R}_{q}$ and the radar operates under a small squint angle [5]. In LAWNs, $vT_{a}$ of the UAV can reach on the order of hundreds of meters, while its altitude can be on the order of several thousands of meters. In addition, the small-squint-angle condition is naturally satisfied in broadside mode. Therefore, the approximation in (8) holds.
Then substituting (8) into (7) and rearranging terms yields
@@
where $\alpha_{q}=\bar{\alpha}_{q}e^{-j\frac{4\pi}{c}f_{c}\bar{R}_{q}}$. In addition to separated range and azimuth terms, the range-cell migration (RCM) term makes conventional 2D-FFT approach in [9, 18] difficult to achieve independent range-Doppler focusing. While the CP effectively mitigates ISI, the high-mobility trajectory of the UAV introduces a time-variant phase coupling across the TF grid. This RCM effect is essentially the manifestation of time-variant frequency selectivity, where the channel response evolves across both subcarriers and symbols. The subsequent TF-domain filtering framework is designed to embrace these dynamics by regularizing the random waveform, thereby restoring the coherence required for high-resolution imaging.
@@
$\mathbf{Y}=\mathbf{H}\odot\mathbf{S}+\mathbf{Z},$ (10)
where $\mathbf{S}\in\mathbb{C}^{N\times M}$ and $\mathbf{Z}\in\mathbb{C}^{N\times M}$ represent the random symbol and noise matrices, respectively, and $\mathbf{H}\in\mathbb{C}^{N\times M}$ denotes the imaging channel matrix, expressed as
$\mathbf{H}=\sum\nolimits^{Q}_{q=1}\alpha_{q}\underbrace{\mathbf{b}\mathbf{c}^{H}\odot\mathbf{E}}_{\mathbf{H}_{q}}.$ (11)
Here, $\mathbf{b}\in\mathbb{C}^{N\times 1}$ with the $n$th element as $\mathbf{b}]_{n}=e^{-j\frac{4\pi}{c}n\Delta f\bar{R}_{q}}$, $\mathbf{c}\in\mathbb{C}^{M\times 1}$ with the $m$th element as $\mathbf{c}]_{m}=e^{-j\frac{4\pi}{c}f_{c}\Delta R_{q,m}}$, and $\mathbf{E}\in\mathbb{C}^{N\times M}$ with the $(n,m)$th element as $\mathbf{E}]_{n,m}=e^{-j\frac{4\pi}{c}n\Delta f\Delta R_{q,m}}$.
@@
The reconstruction of OFDM-SAR imaging can be expressed as a linear model by reformulating (10) as
${\mathbf{y}}=\tilde{\mathbf{H}}\bm{\alpha}+{\mathbf{z}},$ (12)
where ${\mathbf{y}}=\mathrm{vec}(\mathbf{Y})$, $\tilde{\mathbf{H}}=\left[\mathrm{vec}(\mathbf{H}_{1}\odot\mathbf{S}),\cdots,\mathrm{vec}(\mathbf{H}_{Q}\odot\mathbf{S})\right]$, $\bm{\alpha}=[\alpha_{1},\cdots,\alpha_{Q}]^{T}$, and $\mathbf{z}=\mathrm{vec}(\mathbf{Z})$. Since $\mathbf{z}$ is Gaussian and satisfies $\mathbb{E}\{\mathbf{z}\}=\mathbf{0}$ and $\mathbf{R}_{z}=\mathbb{E}\{\mathbf{z}\mathbf{z}^{H}\}=\sigma^{2}\mathbf{I}$, from a parameter estimation perspective, SAR imaging can thus be interpreted as the least-squared (LS) estimation of $\bm{\alpha}$ [45, 44]
$\hat{\bm{\alpha}}=\left(\tilde{\mathbf{H}}^{H}\tilde{\mathbf{H}}\right)^{-1}\tilde{\mathbf{H}}^{H}\mathbf{y}.$ (13)
Since the number of pixels $Q$ is typically enormous, $\tilde{\mathbf{H}}\in\mathbb{C}^{NM\times Q}$ becomes too large to store, making direct LS inversion computationally intractable. Instead, SAR commonly resorts to FFT-based methods, e.g., range-Doppler or Omega-K algorithms [5], which are typically implemented via coherent phase-history processing, such as range compression, azimuth focusing, and Doppler-domain processing. Thanks to the inherent Fourier structure of $\tilde{\mathbf{H}}$ to avoid the direct inversion of a prohibitively large matrix, these widely used algorithms can be interpreted as computationally efficient approximations to the LS solution, which achieve near-optimal focusing performance with significantly reduced complexity. This connection provides a useful bridge between the proposed OFDM-based formulation and conventional SAR imaging mechanisms, which is further clarified in Sec. III.
@@ III SAR Imaging Approach
In this section, we elaborate on the procedure of integrating SAR imaging into UAV data backhaul, where the technical roadmap is shown in Fig. 2. There are two key differences compared to the conventional RD algorithm. First, we propose to exploit several TF-domain filtering schemes to allieviate the impact of data randomness. In this work, the TF-domain filtering is not introduced as a new estimator, but as a means to establish the mapping from random communication data symbols to SAR imaging performance. This mapping enables a systematic analysis of how signal randomness propagates through the imaging chain and affects the final image quality. Second, the azimuth signal is not a strict quadratic-phase signal, since the random data imposes amplitude modulation on it, and the approximation of its Doppler spectrum requires further discussion.
Figure 2: Technical roadmap.
@@ III-A TF Filtering
Different from conventional SAR imaging based on chirp waveforms, SAR using OFDM communication signals may suffer from severe imaging degradation due to the modulation randomness and the employed TF-domain filtering schemes. To mitigate this effect, the first step is to compensate for the discrete symbols drawn from a given QAM constellation.
@@
$\mathbf{y}=\mathbf{A}\mathbf{h}+\mathbf{z},$ (14)
where $\mathbf{A}=\mathrm{diag}(\mathrm{vec}(\mathbf{S}))$, and $\mathbf{h}=\mathrm{vec}(\mathbf{H})$ which involves $\bm{\alpha}$. Then we arrive at $\mathbb{E}\{\mathbf{h}\}=\mathbf{0}$ and $\mathbf{R}_{h}=\mathbb{E}\{\mathbf{h}\mathbf{h}^{H}\}=\sum^{Q}_{q=1}\sigma^{2}_{\alpha_{q}}\mathbf{I}$, which can be similarly proved by referring to [18, Lemma 1].
@@
By rewriting the model as (14), we shift the focus to the channel (observation) space, where $\mathbf{h}$ represents the equivalent imaging channel. The diagonal structure of $\mathbf{A}$ decouples the observations, enabling efficient $NM$ parallel estimation. Notably, from a parameter-space perspective like that in (12), estimating $\mathbf{h}$ implicitly contains information of $\bm{\alpha}$. Consequently, this formulation provides a natural two-stage estimation strategy: 1) channel-space estimation: estimate $\mathbf{h}$ independently for each subchannel; 2) parameter-space inversion: recover the physical scatterer parameters $\bm{\alpha}$ from $\mathbf{h}$ in accordance with their linear relationship as $\mathbf{h}=\mathrm{vec}(\mathbf{H})=\mathrm{vec}\left(\sum_{q}\alpha_{q}\mathbf{H}_{q}\right)=\sum_{q}\alpha_{q}\mathrm{vec}\left(\mathbf{H}_{q}\right)$, through 2D-FFT-based algorithms.
Our current aim is to estimate $\mathbf{h}$, formalized as $\hat{\mathbf{h}}=\mathbf{G}\mathbf{y}$, where $\mathbf{G}$ represents the TF-domain filtering matrix. We establish a unified TF-domain estimation framework for OFDM-SAR imaging under random data modulation, where TF-domain filtering is interpreted as estimating the underlying imaging channel rather than merely performing symbol compensation. Unlike conventional OFDM-SAR systems that employ deterministic and periodically repeated waveforms across pulses, the considered framework utilizes OFDM communication signals, where data symbols vary randomly, introducing random modulation effects into the imaging process.
Within this framework, classical strategies such as reciprocal filtering (RF), matched filtering (MF), and Wiener filtering (WF), are revisited as estimators under different optimality criteria. This unified perspective enables a systematic analysis of how different filtering strategies affect imaging quality, and provides new insights into the mapping mechanism between signal randomness and SAR imaging, which is not captured in conventional systems with deterministic waveforms.
@@ III-A1 Reciprocal Filtering (RF)
$\hat{\mathbf{h}}_{\mathrm{RF}}=\arg\min_{\mathbf{h}}\ \|\mathbf{y}-\mathbf{A}\mathbf{h}\|^{2}_{F},$ (15)
@@
$\hat{\mathbf{h}}_{\mathrm{RF}}=\underbrace{\left(\mathbf{A}^{H}\mathbf{A}\right)^{-1}\mathbf{A}^{H}}_{\mathbf{G}_{\mathrm{RF}}}\mathbf{y}.$ (16)
Thanks to the diagonal structure of $\mathbf{A}$, (16) is equivalent to
$\hat{\mathbf{H}}_{\mathrm{RF}}=\frac{\mathbf{Y}\odot\mathbf{S}^{*}}{|\mathbf{S}|^{2}}=\frac{\mathbf{Y}}{\mathbf{S}},$ (17)
where $\hat{\mathbf{H}}_{\mathrm{RF}}$ is the RF estimation of $\mathbf{H}$.
@@ III-A2 Matched Filtering (MF)
MF is designed based on the rule of maximum SNR output, which may be quantified as
$\mathrm{SNR}_{\mathrm{out}}=$ $\frac{\mathbb{E}\{\|\mathbf{G}_{\mathrm{MF}}\mathbf{A}\mathbf{h}\|^{2}_{F}\}}{\mathbb{E}\{\|\mathbf{G}_{\mathrm{MF}}\mathbf{z}\|^{2}_{F}\}}=\mathrm{SNR}_{\mathrm{in}}\cdot\frac{tr\left(\mathbf{G}_{\mathrm{MF}}\mathbf{A}\mathbf{A}^{H}\mathbf{G}^{H}_{\mathrm{MF}}\right)}{tr\left(\mathbf{G}_{\mathrm{MF}}\mathbf{G}^{H}_{\mathrm{MF}}\right)}$ (18)
$\leq$ $\mathrm{SNR}_{\mathrm{in}}\cdot\frac{tr\left(\mathbf{A}\mathbf{A}^{H}\right)tr\left(\mathbf{G}^{H}_{\mathrm{MF}}\mathbf{G}_{\mathrm{MF}}\right)}{tr\left(\mathbf{G}_{\mathrm{MF}}\mathbf{G}^{H}_{\mathrm{MF}}\right)}$
$=$ $\mathrm{SNR}_{\mathrm{in}}\cdot tr\left(\mathbf{A}\mathbf{A}^{H}\right)=NM\cdot\mathrm{SNR}_{\mathrm{in}},$
where $\mathrm{SNR}_{\mathrm{in}}=\sum^{Q}_{q=1}\sigma^{2}_{\alpha_{q}}/\sigma^{2}$, and the maximum $\mathrm{SNR}_{\mathrm{out}}$ is attained when $\mathbf{G}_{\mathrm{MF}}=\beta\mathbf{A}^{H}$. Without loss of generality, we use $\beta=1$, leading to
$\hat{\mathbf{h}}_{\mathrm{MF}}=\mathbf{A}^{H}\mathbf{y}.$ (19)
Consequently, $\mathbf{H}$ is estimated through
$\hat{\mathbf{H}}_{\mathrm{MF}}=\mathbf{Y}\odot\mathbf{S}^{*}.$ (20)
@@ III-A3 Wiener Filtering (WF)
${\mathbf{G}}_{\mathrm{WF}}=\arg\min_{\mathbf{G}}\ \mathbb{E}\{\|\mathbf{h}-\mathbf{G}\mathbf{y}\|^{2}_{F}\},$ (21)
@@
$\hat{\mathbf{h}}_{\mathrm{WF}}=\underbrace{\mathbf{R}_{h}\mathbf{A}^{H}\left(\mathbf{A}\mathbf{R}_{h}\mathbf{A}^{H}+\mathbf{R}_{z}\right)^{-1}}_{\mathbf{G}_{\mathrm{WF}}}\mathbf{y}.$ (22)
@@
$\hat{\mathbf{H}}_{\mathrm{WF}}=\frac{\mathbf{Y}\odot\mathbf{S}^{*}}{|\mathbf{S}|^{2}+\mathrm{SNR}^{-1}_{\mathrm{in}}}.$ (23)
In practice, the prior information $\mathrm{SNR}_{\mathrm{in}}$ for WF can be estimated from the received data, e.g., via background noise estimation and signal power measurement. The proposed analysis assumes that such estimates are available.
Combining (17), (20) and (23), we summarize the TF-domain filtering matrix $\mathbf{G}$ under different schemes as
$\mathbf{G}=\left\{\begin{array}[]{ll}\frac{1}{\mathbf{S}},&\mathrm{RF},\\ \mathbf{S}^{*},&\mathrm{MF},\\ \frac{\mathbf{S}^{*}}{\left|\mathbf{S}\right|^{2}+\mathrm{SNR}_{\mathrm{in}}^{-1}},&\mathrm{WF},\end{array}\right$. (24)
@@
$\mathbf{Y}^{\mathrm{tf}}\overset{\triangle}{=}\hat{\mathbf{H}}=\mathbf{Y}\odot\mathbf{G}=\mathbf{H}\odot\mathbf{S}\odot\mathbf{G}+\mathbf{Z}\odot\mathbf{G},$ (26)
@@
$\mathrm{MSE}=$ $\mathbb{E}\left\{\left\|\hat{\mathbf{H}}-\mathbf{H}\right\|^{2}_{F}\right\}$ (28)
$=$ $\mathbb{E}\left\{\left\|\mathbf{H}\odot\mathbf{S}\odot\mathbf{G}-\mathbf{H}+\mathbf{Z}\odot\mathbf{G}\right\|^{2}_{F}\right\}$
@@
where the third equality holds due to the facts that: 1) the statistical independence between $\mathbf{H}$ and $\mathbf{Z}$, and 2) the statistical characteristics of random variables in (28) are invariant across subcarriers and symbols. Readers are referred to [9] for a detailed derivation.